Find the formula for a given sequence

3/6, 8/25, 15/62, 24/123, 35/214


1
Expert's answer
2021-02-24T07:47:04-0500

f(n) = (n*(n+2))/[n*(n+2)*(n+1) +(n-1)], where f(1) = 3/6